# Efficient Beam Selection for ISAC in Cell-Free Massive MIMO via Digital Twin-Assisted Deep Reinforcement Learning

### v1

@@ I Introduction
Integrated sensing and communication (ISAC) represents a key advancement for future sixth-generation (6G) networks, facilitating the simultaneous operation of communication and radar sensing within the same spectral resources [1]. This dual functionality not only supports high-data-rate applications but also enhances spectral efficiency [2, 3]. ISAC systems within cellular networks have attracted significant attention, particularly in beamforming design [4, 5], parameter estimation [6], and resource allocation [7], typically involving a single base station (BS), where the sensing transmitter and receiver are co-located and operate in full-duplex mode [8]. In contrast, multi-static sensing leverages multiple spatially distributed receivers, offering diversity gains without necessitating full-duplex operation [9]. Therefore, the integration of ISAC with cell-free systems gains increasing attention [10, 11], in which a large number of distributed access points (APs) collaborate to provide the communication and sensing services without traditional cell boundaries [6]. Through backhaul links, a central processing unit (CPU) connects all APs, enabling cooperative communication that enhances coverage and spectral efficiency [12]. Cell-free ISAC addresses the key limitations of single-AP ISAC systems, such as restricted observation angles that hinder target detection, by exploiting diverse sensing observations across multiple receivers, significantly enhancing both communication and sensing performance [13].
@@
As a model-free algorithm, DRL optimizes decisions through agent-environment interaction, rather than conducting global modeling of the environment [19]. It does not rely on precise environmental parameters; instead, it is trained by iteratively updating an online policy according to the environment feedback. To be specific, authors in [20] developed a DRL-based beam training method, interacting with the environment to capture dynamic spatial patterns and decide the beam search range without precise CSI. Authors in [21] introduced a DRL framework to jointly optimize the beam selection and digital precoding matrices, subject to constraints on transmit power and the selection matrix structure. To achieve a dynamic trade-off between beam selection quality and beam sweep overhead, authors in [22] proposed a novel action space in which actions can dynamically adjust the size of the beam sweep subset based on high- and low-frequency channel propagation laws. By formulating a cognitive multi-target detection problem, authors in [23] optimized transmitted waveforms to ensure satisfactory detection performance under pre-determined false alarm rate constraint. However, in certain scenarios, such as autonomous driving, generating online data is impractical due to high costs and associated risks [24]. Thus, the contradiction between online DRL algorithms and practical deployment is one of the biggest obstacles to their widespread adoption. Fortunately, offline DRL was proposed as a promising solution to address this issue [25]. It mitigates the high costs of online interaction by utilizing only previously collected data, which typically contains sequences of state-action transitions, to learn the optimal policy without exploring or interacting with the environment during the learning process. Nevertheless, since agents rely on a fixed dataset, the distribution of states and actions within the dataset often fails to cover the full range of potential situations the agent may encounter during deployment. This results in a distribution shift [26], where the learned policy may overestimate or underperform in unexplored regions of the state space, causing the agent to perform poorly in real-world applications. Although the conservative Q-learning (CQL) framework proposed in [27] prevented overly optimistic estimate and reduced policy deviation from the dataset by constraining the Q-value, its conservativeness may hinder policy improvement and limit the exploration of better actions. To enhance the diversity and realism of the offline dataset across various scenarios, we resort to digital twin (DT) technology.
@@ II-A Signal Model
Let $s_{c,k}(t_{s})$ denote the communication signal for the $k$-th UE and $s_{r}(t_{s})$ denote the radar waveform for target detection at time instant $t_{s}\in\mathcal{T}_{s}$, where $\mathcal{T}_{s}\buildrel\textstyle.\over{=}[0,T_{s}]$ represents the duration of a block with a total duration time of $T_{s}$. The ISAC signal transmitted by the transmit AP is given by
@@
where $\mathbf{f}_{k}$ is the beamforming vector designed for the $k$-th UE, and $\mathbf{w}\in\mathbb{C}^{M\times 1}$ is the sensing beamforming vector.
@@
For ease of representation, we define $\gamma_{n}\buildrel\textstyle.\over{=}\alpha_{n}\sqrt{\beta_{n}}$ and $\mathbf{H}_{n}\buildrel\textstyle.\over{=}\mathbf{a}\left(\phi_{n}\right)\mathbf{a}^{\mathrm{T}}\left(\phi_{0}\right)$ as the unknown reflection coefficient and the end-to-end target response matrix from the transmit AP to the target to the $n$-th receive AP, respectively. The received echo signal in (2) can be rewritten as
@@
Without loss of generality, we process the received signal by a matched filter (MF) $s(t_{s})$ tuned at the estimated delay $\hat{\tau}_{n}$, and assume that the MF can perfectly tuned to the target delay [36, 23], which means $\hat{\tau}_{n}=\tau_{n}$. Consequently, we have
[…]
@@ II-B Joint Target Detection
where $\boldsymbol{\eta}_{n}\buildrel\textstyle.\over{=}\mathbf{H}_{n}\mathbf{w}$ is defined as the reflected signal vectors for the $n$-th receive AP. The null hypothesis $\mathcal{H}_{0}$ represents the case that the sensing area contains only disturbance, which encompasses background clutter and noise, with no actual target present. Conversely, the alternative hypothesis $\mathcal{H}_{1}$ asserts the presence of a target within the sensing area in addition to any existing clutter and noise.
@@
$c_{0_{n}}\buildrel\textstyle.\over{=}\|\mathbf{y}_{n}\|^{2}.$ (8)
@@
$c_{1_{n}}(\gamma_{n})\buildrel\textstyle.\over{=}\|\mathbf{y}_{n}-\gamma_{n}\boldsymbol{\eta}_{n}\|^{2}.$ (10)
@@
$\bar{\Lambda}\left(\mathbf{y}_{n}\right)=\frac{f_{1}(\{\mathbf{y}_{n}\}_{n};\gamma_{n}|\mathcal{H}_{1})}{f_{0}(\{\mathbf{y}_{n}\}_{n}|\mathcal{H}_{0})}\stackrel{{\scriptstyle\mathcal{H}_{1}}}{{\underset{\mathcal{H}_{0}}{\gtrless}}}\bar{\lambda},$ (14)
@@
$\Lambda(\mathbf{y}_{n})$ =.$\buildrel\textstyle.\over{=}\ln\bar{\Lambda}(\mathbf{y}_{n})$ (15)
@@
where $\lambda\buildrel\textstyle.\over{=}\ln\bar{\lambda}$.
@@
where $\chi^{2}_{2}$ is a chi-squared random variable with $2$ degrees of freedom. Since $\Lambda(\mathbf{y}_{n})$ is the sum of $N$ chi-squared random variables, we have
$\Lambda(\mathbf{y}_{n})\sim\chi^{2}_{2N},$ (17)
@@
[…]
where $p_{\Lambda\mid\mathcal{H}_{0}}$ is the PDF of $\Lambda\left(\mathbf{y}_{n}\right)$ under $\mathcal{H}_{0}$ and $P_{\mathrm{FA}}$ is defined as the false alarm rate. For the chi-squared distribution,
@@
In this case, $\Lambda(\mathbf{y}_{n})$ is a weighted sum of non-central chi-squared variables and obeys
$\Lambda(\mathbf{y}_{n})\sim\sum_{n}\chi^{2}_{2}(\lambda^{\mathrm{nc}}_{n}),$ (22)
@@
$\Lambda(\mathbf{y}_{n})\sim\chi^{2}_{2N}\left(\sum\nolimits_{n}\lambda^{\mathrm{nc}}_{n}\right).$ (23)
@@
[…]
where $p_{\Lambda\mid\mathcal{H}_{1}}$ is the PDF of $\Lambda\left(\mathbf{y}_{n}\right)$ under $\mathcal{H}_{1}$.
@@ II-C Problem Formulation
III Online DRL-Based Beam Tracking Method
@@ III-A1 State
where $\operatorname{sgn}\left(\cdot\right)$ is the sign function, $a_{t}$ is the action operated at $t$-th stage. Consequently, the state vector at $t$-th stage is mathematically defined as $\mathbf{s}_{t}\buildrel\textstyle.\over{=}[\mu_{t},{P}_{D}^{t},f_{g}^{t}]^{\mathrm{T}}$.
@@ III-A3 Reward
Since our goal is to minimize the total number of tracking steps for the agent accomplishing the target sensing task, we consider the term ‘$-1$’ as the reward to penalize the agent to finish the task as soon as possible. Maximizing the cumulative reward is equivalent to minimizing the tracking steps. Therefore, the reward function is designed as follows:
@@ III-B Reward Shaping
As illustrated by the CDF curve in Fig. 2, most beam codewords fail to detect the target, with only a few satisfying the detection performance criteria. Thus, the designed reward function in (27) closely aligns with the optimization objective, however, is severely sparse, exhibiting a flat landscape at $-1$ with a single peak at $b_{0}$ when the goal is achieved. Such a situation will require the agent to spend a significant amount of time on exploration.
@@
is designed only related to the current state. Sigmoid function is employed to emphasize slight variations in low values of ${P}_{D}(t)$.
@@
III-C Online Deep Reinforcement Learning-Based Beamforming
@@
$Q(\mathbf{s}_{t},a_{t};\theta_{q})\buildrel\textstyle.\over{=}\mathbb{E}\left[\sum_{t^{\prime}=0}^{\infty}\rho^{t^{\prime}}r_{t^{\prime}}|\mathbf{s}_{0}=\mathbf{s}_{t},a_{0}=a_{t};\theta_{q}\right],$ (30)
@@ III-C1 Q-Network Structure
$\bar{P}_{D}^{t}=\sigma_{v}(P_{D}^{t})\buildrel\textstyle.\over{=}\frac{\frac{1}{1+e^{-b_{3}\cdot{P}_{D}^{t}}}-\frac{1}{2}}{\frac{1}{1+e^{-b_{3}}}-\frac{1}{2}}.$ (32)
This design guarantees a continuous mapping of values within $[0,1]$ to the same range, while significantly enhancing the sensitivity of the network to capture subtle variations in ${P}_{D}^{t}$ around 00.
@@ III-C2 Interactive Process
However, such trial-and-error method often requires a lot of interaction costs. As a result, we introduce offline DRL and enhance the training efficiency by employing the digital twin technique.

### v2

@@ I Introduction
Integrated sensing and communication (ISAC) represents a key advancement for future sixth-generation (6G) networks, facilitating the simultaneous operation of communication and radar sensing within the same spectral resources [21, 43]. This dual functionality not only supports high-data-rate applications but also enhances spectral efficiency [38, 9]. ISAC systems within cellular networks have attracted significant attention, particularly in beamforming design [14, 20], parameter estimation [34], and resource allocation [16, 44], typically involving a single base station (BS), where the sensing transmitter and receiver are co-located and operate in full-duplex mode [5]. In contrast, multi-static sensing leverages multiple spatially distributed receivers, offering diversity gains without necessitating full-duplex operation [3]. Therefore, the integration of ISAC with cell-free systems gains increasing attention [22, 6], in which a large number of distributed access points (APs) collaborate to provide the communication and sensing services without traditional cell boundaries [34]. Through backhaul links, a central processing unit (CPU) connects all APs, enabling cooperative communication that enhances coverage and spectral efficiency [24]. Cell-free ISAC addresses the key limitations of single-AP ISAC systems, such as restricted observation angles that hinder target detection, by exploiting diverse sensing observations across multiple receivers, significantly enhancing both communication and sensing performance [37].
@@
As a model-free algorithm, DRL optimizes decisions through agent-environment interaction, rather than conducting global modeling of the environment [35]. It does not rely on precise environmental parameters; instead, it is trained by iteratively updating an online policy according to the environment feedback. To be specific, authors in [39] developed a DRL-based beam training method, interacting with the environment to capture dynamic spatial patterns and decide the beam search range without precise CSI. Authors in [15] introduced a DRL framework to jointly optimize the beam selection and digital precoding matrices, subject to constraints on transmit power and the selection matrix structure. To achieve a dynamic trade-off between beam selection quality and beam sweep overhead, authors in [46] proposed a novel action space in which actions can dynamically adjust the size of the beam sweep subset based on high- and low-frequency channel propagation laws. By formulating a cognitive multi-target detection problem, authors in [1] optimized transmitted waveforms to ensure satisfactory detection performance under pre-determined false alarm rate constraint. Additionally, several recent studies have successfully combined ISAC and DRL to address challenges in trajectory planning and resource allocation under dynamic aerial environments [36]. However, in many practical scenarios, generating online data through continuous environment interaction is infeasible due to the high operational cost, safety concerns, and potential risks involved [19]. Thus, the contradiction between online DRL algorithms and the practical requirements of low latency and low energy consumption poses one of the major obstacles to their widespread adoption in future 6G networks. Fortunately, offline DRL was proposed as a promising solution to address this issue [10]. It mitigates the high costs of online interaction by utilizing only previously collected data, which typically contains sequences of state-action transitions, to learn the optimal policy without exploring or interacting with the environment during the learning process. Nevertheless, since agents rely on a fixed dataset, the distribution of states and actions within the dataset often fails to cover the full range of potential situations the agent may encounter during deployment. This results in a distribution shift [17], where the learned policy may overestimate or underperform in unexplored regions of the state space, causing the agent to perform poorly in real-world applications. Although the conservative Q-learning (CQL) framework proposed in [18] prevented overly optimistic estimate and reduced policy deviation from the dataset by constraining the Q-value, its conservativeness may hinder policy improvement and limit the exploration of better actions. To enhance the diversity and realism of the offline dataset across various scenarios, we resort to digital twin (DT) technology.
@@
The remainder of this paper is organized as follows. Section II presents the system model and formulates the beamforming problem. Section III outlines the online DRL method as a baseline. Section IV introduces the proposed digital twin-assisted offline DRL framework and elaborates on the design of the interactive beamforming procedure. Section V provides numerical results and discussions, and Section VI concludes the paper.
@@ II-A Signal Model
Let $s_{c,k}(t_{s})$ denote the communication signal for the $k$-th UE and $s_{r}(t_{s})$ denote the radar waveform for target detection at time instant $t_{s}\in\mathcal{T}_{s}$, where $\mathcal{T}_{s}\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}[0,T_{s}]$ represents the duration of a block with a total duration time of $T_{s}$. The ISAC signal transmitted by the transmit AP is given by
@@
where $\mathbf{f}_{k}\in\mathbb{C}^{M\times 1}$ is the beamforming vector designed for the $k$-th UE, and $\mathbf{w}\in\mathbb{C}^{M\times 1}$ is the sensing beamforming vector.
@@
For ease of representation, we define $\gamma_{n}\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\alpha_{n}\sqrt{\beta_{n}}$ and $\mathbf{H}_{n}\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\mathbf{a}\left(\phi_{n}\right)\mathbf{a}^{\mathrm{T}}\left(\phi_{0}\right)\in\mathbb{C}^{M\times M}$ as the unknown reflection coefficient and the end-to-end target response matrix from the transmit AP to the target to the $n$-th receive AP, respectively. The received echo signal in (2) can be rewritten as
@@
Without loss of generality, we process the received signal by a matched filter (MF) $s(t_{s})$ through correlation with the known transmit waveform, where the peak of the correlation output provides an estimate $\hat{\tau}_{n}$ of the target delay. The MF is optimal for maximizing the output SNR with a known transmit waveform, which directly governs the detection probability. In the absence of external interference, using the MF introduces no intrinsic performance loss. For analytical tractability, we adopt the ideal assumption that the MF can perfectly tuned to the target delay [11, 1], which means $\hat{\tau}_{n}=\tau_{n}$. Consequently, we have
$\mathbf{y}_{n}=$ $\frac{1}{T_{s}}\int_{\mathcal{T}_{s}}\mathbf{y}_{n}(t_{s})s_{r}^{*}(t_{s}-\hat{\tau}_{n})dt_{s}=\gamma_{n}\mathbf{H}_{n}\mathbf{w}+\bar{\boldsymbol{\epsilon}}_{n},$ (5)
where $\frac{1}{T_{s}}\int_{\mathcal{T}_{s}}s_{c,k}(t_{s})s_{r}^{*}(t_{s})dt_{s}=0$ is assumed, and $\bar{\boldsymbol{\epsilon}}_{n}=\frac{1}{T_{s}}\int_{\mathcal{T}_{s}}\boldsymbol{\epsilon}_{n}(t)s_{r}^{*}(t_{s}-\hat{\tau}_{n})dt_{s}$ denotes the equivalent noise after MF processing. By aggregating all the signals $\mathbf{y}_{n},n=1,\ldots,N$, the CPU conducts cooperative target detection based on the composite information derived from each receive AP.
@@ II-B Joint Target Detection
where $\boldsymbol{\eta}_{n}\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\mathbf{H}_{n}\mathbf{w}$ is defined as the reflected signal vectors for the $n$-th receive AP. The null hypothesis $\mathcal{H}_{0}$ represents the case that the sensing area contains only disturbance, which encompasses background clutter and noise, with no actual target present. Conversely, the alternative hypothesis $\mathcal{H}_{1}$ asserts the presence of a target within the sensing area in addition to any existing clutter and noise.
@@
$c_{0_{n}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\|\mathbf{y}_{n}\|^{2}.$ (8)
@@
$c_{1_{n}}(\gamma_{n})\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\|\mathbf{y}_{n}-\gamma_{n}\boldsymbol{\eta}_{n}\|^{2}.$ (10)
@@
$\bar{\Lambda}\left(\mathbf{y}\right)=\frac{f_{1}(\{\mathbf{y}_{n}\}_{n};\gamma_{n}|\mathcal{H}_{1})}{f_{0}(\{\mathbf{y}_{n}\}_{n}|\mathcal{H}_{0})}\stackrel{{\scriptstyle\mathcal{H}_{1}}}{{\underset{\mathcal{H}_{0}}{\gtrless}}}\bar{\lambda},$ (14)
@@
$\Lambda(\mathbf{y})$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\ln\bar{\Lambda}(\mathbf{y})$ (15)
@@
where $\lambda\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\ln\bar{\lambda}$.
@@
where $\chi^{2}_{2}$ is a chi-squared random variable with $2$ degrees of freedom. Since $\Lambda(\mathbf{y})$ is the sum of $N$ chi-squared random variables, we have
$\Lambda(\mathbf{y})\sim\chi^{2}_{2N},$ (17)
@@
$\operatorname{Pr}\left\{\Lambda(\mathbf{y})>\lambda\mid\mathcal{H}_{0}\right\}=\int_{\lambda}^{\infty}p_{\Lambda\mid\mathcal{H}_{0}}\left(a\mid\mathcal{H}_{0}\right)da\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}P_{\mathrm{FA}},$ (18)
where $p_{\Lambda\mid\mathcal{H}_{0}}$ is the PDF of $\Lambda\left(\mathbf{y}\right)$ under $\mathcal{H}_{0}$ and $P_{\mathrm{FA}}$ is defined as the false alarm rate. For the chi-squared distribution,
@@
In this case, $\Lambda(\mathbf{y})$ is a weighted sum of non-central chi-squared variables and obeys
$\Lambda(\mathbf{y})\sim\sum_{n}\chi^{2}_{2}(\lambda^{\mathrm{nc}}_{n}),$ (22)
@@
$\Lambda(\mathbf{y})\sim\chi^{2}_{2N}\left(\sum\nolimits_{n}\lambda^{\mathrm{nc}}_{n}\right).$ (23)
@@
$\operatorname{Pr}\left\{\Lambda(\mathbf{y})>\lambda\mid\mathcal{H}_{1}\right\}=\int_{\lambda}^{\infty}p_{\Lambda\mid\mathcal{H}_{1}}\left(a\mid\mathcal{H}_{1}\right)da\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}P_{D},$ (24)
where $p_{\Lambda\mid\mathcal{H}_{1}}$ is the PDF of $\Lambda\left(\mathbf{y}\right)$ under $\mathcal{H}_{1}$.
@@
Remark 2.
While Eq. (6) assumes i.i.d. AWGN across APs, this assumption is mainly adopted for analytical tractability. In practice, the noises at spatially separated APs are independent but may exhibit distinct variances due to hardware or environmental differences. In such cases, the detection statistic becomes a sum of independent but non-identically distributed non-central chi-squared variables, whose characteristic function is
$\varphi_{\Lambda\mid\mathcal{H}_{1}}(t)=\prod_{n=1}^{N}(1-2j\sigma_{n}^{2}\|\boldsymbol{\eta}_{n}\|^{2}t)^{-1}e^{\frac{j\left|\hat{\gamma}_{n}\|\boldsymbol{\eta}_{n}\|^{2}\right|^{2}\cdot t}{1-2j\sigma_{n}^{2}\|\boldsymbol{\eta}_{n}\|^{2}\cdot t}},$ (25)
and its PDF can be obtained via inverse Fourier transform. This generalization does not affect the independence assumption but slightly modifies the analytical expression of the detection probability.
@@ II-C Problem Formulation
III Online Deep Reinforcement Learning-Based Beam Tracking Method
@@ III-A1 State
where $\operatorname{sgn}\left(\cdot\right)$ is the sign function, $a_{t}$ is the action operated at $t$-th stage. When $f_{g}^{t}=1$, the beam adjustment enhances the detection probability compared with the previous step; when $f_{g}^{t}=-1$, it indicates a degradation; and $f_{g}^{t}=0$ corresponds to the initialization stage. This interpretable variable thus provides a concise and physically meaningful feedback that guides the agent to progressively enhance beam alignment and detection performance. Consequently, the state vector at $t$-th stage is mathematically defined as $\mathbf{s}_{t}\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}[\mu_{t},{P}_{D}^{t},f_{g}^{t}]^{\mathrm{T}}$.
@@ III-A3 Reward
Since our goal is to minimize the total number of tracking steps for the agent accomplishing the target sensing task, we consider the term ‘1-1’ as the reward to penalize the agent to finish the task as soon as possible. Maximizing the cumulative reward is equivalent to minimizing the tracking steps. Therefore, the reward function is designed as follows:
@@ III-B Reward Shaping
As illustrated by the CDF curve in Fig. 2, most beam codewords fail to detect the target, with only a few satisfying the detection performance criteria. Thus, the designed reward function in (28) closely aligns with the optimization objective, however, is severely sparse, exhibiting a flat landscape at 1-1 with a single peak at $b_{0}$ when the goal is achieved. Such a situation will require the agent to spend a significant amount of time on exploration.
@@
is designed only related to the current state, where $b_{1}$ controls the amplitude and scales the magnitude of the shaping reward, while $b_{2}$ determines the steepness of the sigmoid function. The sigmoid form is employed to emphasize slight variations in low values of ${P}_{D}(t)$.
@@
III-C Online DRL-Based Beam tracking
@@
$Q(\mathbf{s}_{t},a_{t};\theta_{q})\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\mathbb{E}\left[\sum_{t^{\prime}=0}^{\infty}\rho^{t^{\prime}}r_{t^{\prime}}|\mathbf{s}_{0}=\mathbf{s}_{t},a_{0}=a_{t};\theta_{q}\right],$ (31)
@@ III-C1 Q-Network Structure
$\bar{P}_{D}^{t}=\sigma_{v}(P_{D}^{t})\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\frac{\frac{1}{1+e^{-b_{3}\cdot{P}_{D}^{t}}}-\frac{1}{2}}{\frac{1}{1+e^{-b_{3}}}-\frac{1}{2}}.$ (33)
This design guarantees a continuous mapping of values within $[0,1]$ to the same range, while significantly enhancing the sensitivity of the network to capture subtle variations in ${P}_{D}^{t}$ around 0.
@@ III-C2 Interactive Process
III-C3 Practical Implementation
During the practical deployment stage, the well-trained Q-network is directly used for real-time decision-making. The resulting policy is characterized by the optimal action-value function $Q^{*}(\mathbf{s},a;\theta_{q}^{*})$, where $\theta_{q}^{*}$ denotes the optimal network parameters obtained after convergence. Given the current observation $\mathbf{s}_{t}$, the transmitter selects its beam index by
$a_{t}=\arg\max_{a}Q^{*}(\mathbf{s}_{t},a;\theta_{q}^{*}),$ (37)
requiring only a single forward pass with negligible latency.